OpenStudy (anonymous):

Medal Two six-sided dice are rolled. Find the probability of the given outcome (express your answers as fractions in lowest terms). (a) The sum of the dots on the upward faces is 11. (b) The sum of the dots on the upward faces is greater than 5.

OpenStudy (ahsome):

(A) Think of the possibilities that could lead to 11: \(5+6=11\) \(6+5=11\) There are 2 possibilities. There are two dices with 6 sides. To find the total possibilities, simply multiply the numbers together. \(6*6=36\) So there are 2 posibilities we want, out of 36. This gives: \[\frac{2}{36}\]\[=\frac{1}{16}\] We could also find the percentage \[Percentage=\frac{1}{16}*100\] \[Percentage=6.25%\]

OpenStudy (anonymous):

so part a is 1/16 in lowest terms

OpenStudy (ahsome):

Yes

jimthompson5910 (jim_thompson5910):

2/36 doesn't reduce to 1/16

jimthompson5910 (jim_thompson5910):

it's close though

OpenStudy (ahsome):

Sorry

OpenStudy (ahsome):

Its \[\frac{1}{18}\]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (ahsome):

Do the same thing for (B) There are a total of 26 combinations that equal to greater than 5 \[\frac{26}{36}\] \[=\frac{13}{18}\]

OpenStudy (anonymous):

13/18 is in lowest terms?

OpenStudy (ahsome):

Combinations for (B) 1+5, 2+4, 3+3, 4+2, 5+1, 1+6, 2+5, 3+4, 4+3, 5+2, 6+1, 2+6, 3+5, 4+4, 5+3, 6+2, 3+6, 4+5, 5+4, 6+3, 4+6, 5+5, 6+4, 6+5. 5+6, 6+6

OpenStudy (ahsome):

@harz360 Yes

jimthompson5910 (jim_thompson5910):

something like this may come in handy http://bestcase.files.wordpress.com/2011/01/dicediagram.jpg

OpenStudy (anonymous):

Here's a visual for the sample space we're working with (where the first row and column are the values of the sides of the two dice and each entry in the table is then the sum of the dice). \[\begin{array}{c|c|c|c|c|c|c} & 1 & 2 & 3 & 4 & 5 & 6\\\hline 1 & 2 & 3 & 4 & 5 & 6 & 7\\\hline 2 & 3 & 4 & 5 & 6 & 7 & 8\\\hline 3 & 4 & 5 & 6 & 7 & 8 & 9\\ \hline 4 & 5 & 6 & 7 & 8 & 9 & 10 \\\hline 5 & 6 & 7 & 8 & 9 & 10 & 11\\\hline 6 & 7 & 8 & 9 & 10 & 11 & 12\end{array}\] In part (a), we notice that the following entries add to 11: \[\begin{array}{c|c|c|c|c|c|c} & 1 & 2 & 3 & 4 & 5 & 6\\\hline 1 & 2 & 3 & 4 & 5 & 6 & 7\\\hline 2 & 3 & 4 & 5 & 6 & 7 & 8\\\hline 3 & 4 & 5 & 6 & 7 & 8 & 9\\ \hline 4 & 5 & 6 & 7 & 8 & 9 & 10 \\\hline 5 & 6 & 7 & 8 & 9 & 10 & \color{red}{11}\\\hline 6 & 7 & 8 & 9 & 10 & \color{red}{11} & 12\end{array}\] There are two such values, and since there are 36 total outcomes here, we now see that the probability that the sum of the two dice is 11 is \(\dfrac{2}{36} = \dfrac{1}{18}\). In part (b), the entries that sum to values larger than 5 are highlighed: \[\begin{array}{c||c|c|c|c|c|c} & 1 & 2 & 3 & 4 & 5 & 6\\\hline\hline 1 & 2 & 3 & 4 & 5 & \color{red}{6} & \color{red}{7}\\\hline 2 & 3 & 4 & 5 & \color{red}{6} & \color{red}{7} & \color{red}{8}\\\hline 3 & 4 & 5 & \color{red}{6} & \color{red}{7} & \color{red}{8} & \color{red}{9}\\ \hline 4 & 5 &\color{red}{6} & \color{red}{7} & \color{red}{8} & \color{red}{9} & \color{red}{10} \\\hline 5 & \color{red}{6} & \color{red}{7} & \color{red}{8} & \color{red}{9} & \color{red}{10} & \color{red}{11}\\\hline 6 & \color{red}{7} & \color{red}{8} & \color{red}{9} & \color{red}{10} & \color{red}{11} & \color{red}{12}\end{array}\] Can you take things from here and find the probability that the sum is larger than 5?

OpenStudy (ahsome):

This could also be useful: http://www.freemathhelp.com/rolling-dice.html

OpenStudy (anonymous):

oh ok thank you so part a is 1/18 and part b is 13/18

OpenStudy (ahsome):

@harz360, Yes :)

OpenStudy (ahsome):

@ChristopherToni, nice diagram :)

OpenStudy (ahsome):

The power of \(LaTeX\) never seems to end

OpenStudy (anonymous):

It was annoying coloring each entry individually. \(\LaTeX\) certainly has it's pros and cons. >_>

OpenStudy (anonymous):

thanks everyone for their help :)

OpenStudy (anonymous):

what would be the median of this 73, 80, 90, 72, 83, 79, 91, 75, 74, 88, 76, and 79 , would it be 73+79/2 = 118.5?

OpenStudy (anonymous):

79+79/2 = 118.5 *

OpenStudy (ahsome):

Yes @ChristopherToni, yes it does. I am pretty happy I can do matrices right now ;) \[\begin{bmatrix} Y& E& A& H \\ \end{bmatrix}\]

jimthompson5910 (jim_thompson5910):

the median is not 118.5

OpenStudy (anonymous):

then 79

jimthompson5910 (jim_thompson5910):

yes because (79+79)/2 = 79

OpenStudy (anonymous):

ok thank you

OpenStudy (ahsome):

First of all, organize the numbers: \[73, 80, 90, 72, 83, 79, 91, 75, 74, 88, 76,79\]\(\text{ORGANIZED}\) \[72, 73, 74, 75, 76, 79, 79, 80, 83, 88, 90, 91\] Use this equation to find the median: \[Median=\frac{n+1}{2}\]\[Median=\frac{12+1}{2}\]\[Median=\frac{13}{2}\]\[Median=6.5th\] Its the 6.5th number. This will be your median: 72, 73, 74, 75, 76, \(79\), \(79\), 80, 83, 88, 90, 91\[Median=\frac{79+79}{2}\]\[Median=79\] Therefore, @harz360, the median is 79

OpenStudy (anonymous):

thank soo much :)

OpenStudy (ahsome):

No problem @harz360 :D